OpenStudy (anonymous):

dertermine the three cubic roots of 8. give the answer in the form a+bj. also give the main root in the exponential form

OpenStudy (phi):

If you put 8 into complex polar form raised to the 1/3 power \[ (8 e^{2\pi i})^{\frac{1}{3}} \]

OpenStudy (phi):

however, you can substitute \( 2 \pi = 4 \pi = 6 \pi \) for the angle and you will find 3 unique roots (before finding repeats)

OpenStudy (amistre64):

i recall noticing that "3"roots would divide the unit circle into 3 equal parts starting at (8+0i) then rotating it 120, and another 120 would be the other roots, if im remembering it right

OpenStudy (amistre64):

 -4+4sqrt(3)i, and -4-4sqrt(3)i .....